# VidCapBench: A Comprehensive Benchmark of Video Captioning for Controllable Text-to-Video Generation

### v1

@@ Abstract
Xinlong Chen1,2thanks: Work done during an internship at Kuaishou Technology., Yuanxing Zhang3, Chongling Rao3, Yushuo Guan3, Jiaheng Liu4, Fuzheng Zhang3, Chengru Song3, Qiang Liu1,2thanks: Corresponding author: qiang.liu@nlpr.ia.ac.cn, Di Zhang3, Tieniu Tan1,2,4, 1New Laboratory of Pattern Recognition (NLPR), Institute of Automation, Chinese Academy of Sciences (CASIA) 2School of Artificial Intelligence, University of Chinese Academy of Sciences 3Kuaishou Technology 4Nanjing University
@@ 1 Introduction
To direct the optimization of T2V models, video captions should be accurate, comprehensive, diverse, concise, and abundant. While caption formats vary across different models Ju et al. (2024); Zheng et al. (2024), the core elements in captions emphasized by T2V models seem to be converging. A practical evaluation of video captions for T2V generation must address three main challenges. Alignment with T2V evaluation: The evaluation should assess whether a video caption adequately covers aesthetics, content, motion, and physical laws, aligning with the key metrics of T2V generation. Diversity and stability: The diversity of evaluation data and the stability of the evaluation approach influence the accurate assessment of caption quality. Impact on T2V generation: The correlation between caption evaluation and T2V performance remains unexplored, lacking evidence on how captions influence the generated videos.
@@ 2 Related Work
Video captioning. The goal of video captioning is to describe a video across several key aspects, aiding understanding Doveh et al. (2023), retrieval Ma et al. (2024), and motion control Wang et al. (2024d). In T2V generation, accurate and detailed video captions can enhance semantic alignment during model training Polyak et al. (2024). Naive captioning models adopt free-form descriptions Chen et al. (2024a); Wang et al. (2024c). To enhance controllability, MiraData Ju et al. (2024), VDC Chai et al. (2024), and Vript Yang et al. (2024b) emphasize specific aspects like subjects, background, and shots, significantly benefiting T2V generation. Other methods describe videos from an event perspective Wang et al. (2024a); He et al. (2024b), capturing temporal information more effectively. Despite advancements in caption controlling Wang et al. (2023); Hua et al. (2024), evaluations with omissions may lead to a seesaw effect where gains in one dimension come at the cost of others, limiting the utility of the captioning model.
@@ 3.1 Preliminaries
Human evaluation. Human evaluation demands annotators to assess captions based on predefined criteria. Experienced annotators deliver accurate and consistent evaluations, along with analysis of erroneous cases, which helps training T2V models. Currently, human annotation primarily employs two methods. In 5-point Likert scale, annotators rate captions on a 5-point scale (1 being worst, 5 being best) based on ground truth. Regarding candidate comparison, pairwise samples are presented to annotators to determine the ratio of good-same-bad and finally conclude with the ELO ratings.
Machine evaluation. Human evaluation can be inconsistent among inexperienced annotators and is generally slower and more expensive. Conversely, automatic machine evaluation of video captions is faster and can provide some guidance for training. Mainstream machine evaluation often utilizes GPT-4 as a judge, which can be divided into two categories:
@@ 3.2 Benchmarking Video Captions
Metrics of caption evaluation. Considering efficiency, cost, and stability, VidCapBench employs a predefined-QA paradigm to evaluate video captioning models. Due to the complexity of video content, answers may involve multiple adjectives, nouns, or verbs. Therefore, the judge model is required to categorize responses into four classes: Wrong ($n_{w}$), indicating explicit mention of the relevant content but with factual errors; Neutral ($n_{n}$), indicating no mention of the relevant content; Partially correct ($n_{p}$), indicating mention of relevant and correct information, but with incomplete descriptions; and Correct ($n_{c}$), indicating complete alignment between the caption and the ground-truth answer. Based on these classifications, we compute the following four metrics to comprehensively assess the performance of captioning models:
@@ 3.4 QA Pairs Split
As discussed in Section 3.2, the difficulty of evaluating the QA pairs has a significant impact on the reliability of machine evaluation. Therefore, we split the total QA pairs based on their evaluation consistency using the same strategy in Section 3.2. In order to achieve a balance between accuracy and efficiency, we identify and segregate QA pairs that fail to receive consistent evaluations within the dimensions of video motion and physical laws, which are more crucial to T2V generation. Consequently, a total of 1,150 QA pairs are classified as VidCapBench-HE, which necessitates human intervention for accurate evaluation. The remaining QA pairs are designated as VidCapBench-AE, which can be evaluated automatically. Figure 3 presents illustrative examples of QA pairs from both categories within VidCapBench. Detailed statistics by dimension are provided in Appendix C.1.
@@ C.1 Detailed Dimensions of Evaluation
C.2 Statistics by Dimension
@@
C.3 Tools in Data Pipeline
@@ Appendix D Discussion
Influence of the judge model. While GPT-4o provides stable judgments as a judge model, its cost and evaluation time pose challenges for iterative T2V development. To investigate the impact of the judge model, we substitute GPT-4o with Qwen2-72B Yang et al. (2024a) as the judge model, using the same evaluation procedure with greedy decoding. The results, summarized in Table 10, reveal an average increase of 40 points in Pre score and a decrease of 35 points in Cov score when Qwen2-72B is employed. Such substantial discrepancies appear implausible and are inconsistent with human evaluations shown in Table 11. Therefore, we caution against the use of locally deployed offline models for the judging process.

### v2

@@ Abstract
Xinlong Chen1,2thanks: Work done during an internship at Kuaishou Technology., Yuanxing Zhang3, Chongling Rao3, Yushuo Guan3, Jiaheng Liu4, Fuzheng Zhang3, Chengru Song3, Qiang Liu1,2thanks: Corresponding author: qiang.liu@nlpr.ia.ac.cn, Di Zhang3, Tieniu Tan1,2,4 1New Laboratory of Pattern Recognition (NLPR), Institute of Automation, Chinese Academy of Sciences (CASIA) 2School of Artificial Intelligence, University of Chinese Academy of Sciences 3Kuaishou Technology 4Nanjing University
@@ 1 Introduction
To direct the optimization of T2V models, video captions should be accurate, comprehensive, diverse, concise, and abundant. While caption formats vary across different models Ju et al. (2024); Zheng et al. (2024), the core elements in captions emphasized by T2V models seem to be converging. A practical evaluation of video captions for T2V generation must address three main challenges:
  • Alignment with T2V evaluation: The evaluation should assess whether a video caption adequately covers aesthetics, content, motion, and physical laws, aligning with the key metrics of T2V generation.
  • Diversity and stability: The diversity of evaluation data and the stability of the evaluation approach influence the accurate assessment of caption quality.
  • Impact on T2V generation: The correlation between caption evaluation and T2V performance remains unexplored, lacking evidence on how captions influence the generated videos.
@@ 2 Related Work
Video captioning. The goal of video captioning is to describe a video across several key aspects, aiding understanding Doveh et al. (2023), retrieval Ma et al. (2024), and motion control Wang et al. (2024d). In T2V generation, accurate and detailed video captions can enhance semantic alignment during model training Polyak et al. (2024). Naive captioning models adopt free-form descriptions Chen et al. (2024a); Wang et al. (2024c). To enhance controllability, MiraData Ju et al. (2024), VDC Chai et al. (2024), and Vript Yang et al. (2024) emphasize specific aspects like subjects, background, and shots, significantly benefiting T2V generation. Other methods describe videos from an event perspective Wang et al. (2024a); He et al. (2024b), capturing temporal information more effectively. Despite advancements in caption controlling Wang et al. (2023); Hua et al. (2024), evaluations with omissions may lead to a seesaw effect where gains in one dimension come at the cost of others, limiting the utility of the captioning model.
@@ 3.1 Preliminaries
Human evaluation. Human evaluation demands annotators to assess captions based on predefined criteria. Experienced annotators deliver accurate and consistent evaluations, along with analysis of erroneous cases, which helps training T2V models. Currently, human annotation primarily employs two methods:
  • 5-point Likert scale: Annotators rate captions on a 5-point scale (1: worst, 5: best) based on ground truth. Each evaluation dimension is assessed independently, with predefined examples illustrating different score levels. To ensure reliability, each example is typically evaluated by a minimum of three annotators, with inter-annotator agreement metrics employed to maintain consistency.
  • Pairwise comparison: Annotators compare two anonymized model outputs for each example, selecting “Caption A is better”, “Caption B is better”, or “Equal quality”. Pairwise comparisons typically utilize the good-same-bad metric, while ranking multiple models can employ ELO scores.
Machine evaluation. Human evaluation can be inconsistent among inexperienced annotators and is generally slower and more expensive. Conversely, automatic machine evaluation is faster and can provide some guidance for training. Mainstream machine evaluation often utilizes GPT-4 as a judge, which can be divided into two categories:
@@ 3.2 Benchmarking Video Captions
Metrics of caption evaluation. Considering efficiency, cost, and stability, VidCapBench employs a predefined-QA paradigm to evaluate video captioning models. Due to the complexity of video content, answers may involve multiple adjectives, nouns, or verbs. Therefore, the judge model is required to categorize responses into four classes:
  • Wrong ($n_{w}$): Responses that explicit mention of the relevant content but with factual errors.
  • Neutral ($n_{n}$): Responses that omit the relevant content entirely.
  • Partially correct ($n_{p}$): Responses that include relevant and correct information, but with incomplete descriptions.
  • Correct ($n_{c}$): Responses that fully align with the ground-truth answer.
Based on these classifications, we compute the following four metrics to comprehensively assess the performance of captioning models:
@@ 3.4 QA Pairs Split
As discussed in Section 3.2, the difficulty of evaluating the QA pairs has a significant impact on the reliability of machine evaluation. Therefore, we split the total QA pairs based on their evaluation consistency using the same strategy in Section 3.2. In order to achieve a balance between accuracy and efficiency, we identify and segregate QA pairs that fail to receive consistent evaluations within the dimensions of video motion and physical laws, which are more crucial to T2V generation. Consequently, a total of 1,150 QA pairs are classified as VidCapBench-HE, which necessitates human intervention for accurate evaluation. The remaining QA pairs are designated as VidCapBench-AE, which can be evaluated automatically. Figure 3 presents illustrative examples of QA pairs from both categories within VidCapBench. Additional examples and detailed statistics of the QA pairs by dimension are provided in Appendix C.
@@ 5 Conclusion
Acknowledgements
This work is jointly sponsored by National Natural Science Foundation of China (62236010, 62141608, 62206291).
@@ C.1 Detailed Dimensions of Evaluation
Figure 15: An example of the QA pairs for a video in VidCapBench.
Figure 16: An example of the QA pairs for a video in VidCapBench.
@@
C.2 Additional QA Pairs Examples
Based on the four dimensions outlined in Section C.1, we have constructed a comprehensive set of QA pairs to assess the quality of captions in the context of T2V generation. In addition to the examples presented in Figure 3, we provide further examples in Figures 15 and 16 to offer a more complete understanding of VidCapBench.
C.3 Statistics by Dimension
@@
C.4 Tools in Data Pipeline
@@ Appendix D Discussion
Influence of the judge model.  While GPT-4o provides stable judgments as a judge model, its cost and evaluation time pose challenges for iterative T2V development. To investigate the impact of the judge model, we substitute GPT-4o with Qwen2-72B and Llama-3.3-70B as the judge model, using the same evaluation procedure with greedy decoding. The results, summarized in Table 10, reveal significant discrepancies in evaluation outcomes. Notably, when Qwen2-72B is employed, we observe an average increase of 40 points in Pre score accompanied by a 35-point decrease in Cov score. Such substantial discrepancies appear implausible and are inconsistent with human evaluations shown in Table 4. Therefore, we caution against the use of locally deployed offline models for the judging process. In contrast, Llama-3.3-70B demonstrates better alignment with GPT-4o, showing comparable Acc, Cov, and Con scores. However, it exhibits systematically higher Pre score that conflict with human judgments in Table 4. Based on these findings, we recommend Llama-3.3-70B as a cost-effective alternative to GPT-4o when budget constraints are a concern, as it provides relatively reliable and meaningful evaluation results. Nevertheless, for more accurate and robust guidance during the development of new models, we still suggest utilizing stronger judge models like GPT-4o whenever feasible.